OpenStudy (anonymous):

What is the length of BC? 7 6 4 2

OpenStudy (amistre64):

british columbia?

OpenStudy (anonymous):

BC is a pretty big place. I live there. I'd guess in the thousands of kilometers range.

OpenStudy (anonymous):



OpenStudy (amistre64):

id try to pythag it first, OR if you have some tools you can just rule it out

OpenStudy (anonymous):

lol! no sorry i have to draw my pictures..

OpenStudy (anonymous):

pythagoras theorem. memorize it. a^2+b^2=c^2 so 8^2+BC^2=10^2 64+?^2=100 ?^2=36 ?=6

OpenStudy (anonymous):

da answer is 6

OpenStudy (amistre64):

one thing to also memorize, is that alot of triangle question use a 3-4-5 setup

OpenStudy (anonymous):

30 60 90 triangle, all triangles have 180 degrees in em

OpenStudy (anonymous):

yeah i guess i have a problem memorizing stuff when it comes to math..but thank you